OpenStudy (anonymous):

Decide whether the pair of lines are paralell, perpendicular, or neither. 3x + 4y=5 4x+3y=9

OpenStudy (anonymous):

solve both of them for Y then you have it like: \[\LARGE y=kx+b\] and for paralel you have: \[\LARGE k_1=k_2\] if prependicular \[\LARGE k_1\cdot k_2=-1 \]

OpenStudy (anonymous):

I am not understanding, so its neither?

OpenStudy (anonymous):

have you solved equations for Y?

OpenStudy (anonymous):

find the slopes of the equations. if they are the same, then they are parallel. and if the product of the slopes is -1, they are perpendicular.

OpenStudy (anonymous):

Is it perpendicular?

OpenStudy (anonymous):

sadly, i have forgotten how to find the slope from the equation. you can google it

OpenStudy (anonymous):

Sorry I wasn't here.. :( I thought you would get the answer... \[\LARGE y=-\frac34x+\frac 53\] and \[\LARGE y=-\frac43x+\frac93\] so you have: \[\LARGE k=-\frac34 \quad \quad \quad ,\quad \quad k_1=-\frac43 \] so we know that \[\LARGE -\frac34\neq -\frac43 \] and that means that \[\LARGE k\neq k_1\] so they're not paralel ... and for prependicular we have: \[\LARGE k\cdot k_1=-1\] \[\LARGE -\frac34\cdot (-\frac43)=+1 \] and we know that \[\LARGE +1\neq -1\] so they're not prependicular either . so they look like: @lovejones2012